OpenStudy (anonymous):

A certain ball of mass M = 0.5 Kg is dropped from a height of 10 meters. This ball will pop if a force greater than 20 Newtons is applied to it . When it hits the ground, it stops completely in .1 seconds. Does the ball pop? What could you do to prevent this?

OpenStudy (shane_b):

The force on impact will equal the change in momentum over the change in time:\[F=\frac{\Delta \rho}{\Delta t}=\frac{mv_f-mv_i}{0.1s}\]You already know the mass...so you need to find the velocity at the time of impact. The simplest way to do this is to use: \[V_f=\pm \sqrt{2gh}=\pm14m/s\]But you can also do it another way - start by calculating the time it takes for the ball to hit the ground from 10m: \[d=V_it+\frac{1}{2}at^2\]\[10m=0m/s^2+\frac{1}{2}(9.8m/s^2)t^2\]\[t=1.429s\] Now determine the ball's velocity at that instant: \[V_f=V_i+at=0m/s+(-9.8m/s^2)(1.429s)=-14m/s\] Note that the sign is only an indication of direction. Now that we know the velocity on impact we can determine the momentum on impact:\[\rho = mv = (0.5kg)(-14m/s)=-7kgm/s\]Of course, the final momentum will be 0 so the change in momentum is 7kgm/s. The force at this instance is:\[F=\frac{\Delta \rho}{\Delta t}=\frac{7kgm/s}{0.1s}=70N\] Therefore, the ball will break. If you want to prevent this you need to increase the amount of time in the above equation so that the total force <= 20N: \[F=20N=\frac{7kgm/s}{Xs}\]Solving for X you get 0.35s. So the impact must occur over a time interval of > 0.35s.

OpenStudy (shane_b):

*That should have been a >= sign in the last sentence.